studybudy123:

Write the first ten terms of a sequence whose first term is -10 and whose common difference is -2.

mhanifa:

Start with -10 and add -2 to each subsequent term: -10, -10 - 2 = -12, -12 - 2 = -14 and so on until you get all 10 terms

Extrinix:

What mhanifa is saying:

-10 -12 -14 etc. -10-2 -12-2 -14-2 etc.